OpenStudy (anonymous):

Could someone help me with this question please, 2. Suppose that f is a function with f(25)=61 and f'(25)=3. If you have no further information, what is your best estimate of f(26.5)?

OpenStudy (blacksteel):

The derivative of a function is the slope of the function at that point. Thus, if f'(25) = 3, then we can estimate that increasing the input value of the function from 25 to 26.5, a change of 1.5, will increase the output of the function by 1.5*3 = 4.5. Thus, given no other information, the best estimate we can give for the value of this function is f(26.5) = 61 + 4.5 = 65.5

OpenStudy (anonymous):

Ohh okay thanks a lot for your help, I appreciate it.